# NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

### v1

@@ 3.2 Neural Network-ResDMD
where $V$ is a matrix in which each column is an eigenvector $\mathbf{v}_{i}$ of Koopman matrix $K$. Thus, with a fixed dictionary function $\mathbf{\Psi}$, the explicit form for the optimal Koopman matrix $\tilde{K}$ can be directly computed as
$\tilde{K}=G^{\dagger}A$ (3.6)
where $G=\frac{1}{m}\Psi_{X}^{*}\Psi_{X},A=\frac{1}{m}\Psi_{X}^{*}\Psi_{Y}$.
Remark.
Typically a regularization term is needed to enhance stability. Here we add a small perturbation, i.e., $\tilde{K}=(G+\sigma I)^{-1}A$ for some small number $\sigma>0$.
As shown in (3.6), NN-ResDMD provides an explicit expression for $\tilde{K}$ given the optimal dictionary function $\mathbf{\Psi}$, allowing for the direct computation of Koopman eigenpairs. The optimization problem in Equation 3.5 is to minimize the error along the eigen-basis, in contrast to the optimization problem $\|\Psi_{Y}-\Psi_{X}K\|_{F}^{2}$ for EDMD, thereby yielding different optimal $\mathbf{\Psi}$ compared to EDMD. Therefore, although the $K$ update procedure appear identical to the EDMD approach, they originate from different theoretical foundations and serve different optimization purposes. Additionally, it automatically optimizes basis functions using neural networks, removing the need for manual selection. Since NN-ResDMD is based on the ResDMD framework, it also retains the theoretical convergence guarantees that EDMD lacks: EDMD has convergence results under strong assumptions, such as requiring the Koopman operator to be bounded (Assumption 2 in [20]), ResDMD requires only that the operator is closed and densely defined.
In NN-ResDMD, neural networks parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$ to minimize the total residual $J(\theta)$, as defined in (3.4). The feedforward neural network generates the dictionary functions based on data snapshots, and the total residual is given by:
$J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2}$ (3.7)
@@
where $K(\theta)$ and $V(\theta)$ depend on $\theta$. The Koopman matrix $\tilde{K}(\theta)$ is computed as:
$\tilde{K}(\theta)=G(\theta)^{\dagger}A(\theta)$ (3.8)
with $G(\theta)=\frac{1}{m}\Psi_{X}(\theta)^{*}\Psi_{X}(\theta)$ and $A(\theta)=\frac{1}{m}\Psi_{X}(\theta)^{*}\Psi_{Y}(\theta)$.
@@
The algorithm alternates between updating $K(\theta)$ via least squares and optimizing $\theta$ using gradient descent until $J(\theta)$ converges, yielding the approximated Koopman spectrum and optimized dictionary functions. While it is possible to optimize both $K(\theta)$ and $J(\theta)$ simultaneously, as done in [21] and [22], our separate procedure ensures computational efficiency and numerical stability compared to the coupled optimization case.
@@
Computing Algorithm In our neural networks implementation, we include some non-trainable basis outputs to enhance the dictionary functions. Specifically, we add a vector of ones and the coordinates of the state space as non-trainable basis in the output layer, which help avoid trivial solutions, i.e., $J=0$ for some initial $\theta$. For the network architecture, we build a three-layer Feedforward Network where each hidden layer size can be specified during training. We use the hyperbolic tangent (tanh) function as the activation function for the hidden layers. In terms of optimization, we employ the Adam optimizer for updating the network parameters. Adam is particularly well-suited for this task due to its ability to adapt the learning rate for each parameter, which can lead to faster convergence in the alternating optimization process between the network parameters and the Koopman matrix. The computing steps are illustrated in the following Algorithm 1.
@@
1 Initialize $\theta$, thus initializing $\mathbf{\Psi}(\theta)$ ;
@@ 6       Update $\tilde{K}(\theta)=(G(\theta)+\sigma I)^{-1}A(\theta)$ and $V(\theta)$ ;
8
Output: $\tilde{K}(\theta)$, eigenpairs $\{(\lambda_{i},\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i})\}_{i=1}^{N_{K}}$ and pseudospectrum $\{z_{j}:\tau_{j}<\varepsilon\}$.
9
@@
While the practical advantages of NN-ResDMD are demonstrated through experiments, it is important to note its computational demands. The algorithm’s computational complexity stems primarily from its iterative optimization process. Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters. Although individual gradient steps are computationally lightweight for standard network architectures, the algorithm’s efficiency issue lies in its repeated least-squares optimizations. Compared to standard single least-squares computation as in most numerical algorithms, NN-ResDMD requires multiple iterations to achieve convergence, with stochastic gradient descent methods showing a theoretical $O(1/n)$ convergence rate. However, the method’s nonlinear optimization nature also presents challenges for establishing concrete convergence bounds and error estimates.
If the continuous spectrum of the Koopman operator is of interest, following the ResDMD paper’s idea, we can scan candidate spectrum values within a grid in the complex plane using the residuals. Specifically, we compute $\tau_{j}=\min_{\mathbf{v}_{i}\in\mathbb{C}^{N_{k}}}\widehat{\textit{res}}(z_{j},\mathbf{\Psi}(\theta)\mathbf{v}_{i})$, where $\tau_{j}$ is the minimum residual for a grid point $z_{j}\in\mathbb{C}$. The approximated whole spectrum containing the continuous spectrum is then given by $\{z_{j}:\tau_{j}<\varepsilon\}$. More details can be found in [12].
While the practical advantages of NN-ResDMD are demonstrated through experiments, it’s also worth noting that the method has theoretical underpinnings [23, 24] that support its convergence properties. A brief discussion on the convergence aspects of NN-ResDMD, leveraging existing results from approximation theory in Barron spaces, is provided in Appendix A.3. This discussion offers insights into how the neural network component of NN-ResDMD contributes to its effectiveness in approximating complex dynamical systems.
@@ 4 Application in physical and biological systems
In this chapter, we present three examples demonstrating the effectiveness of NN-ResDMD in estimating the key quantities of Koopman Mode Decomposition (KMD): spectrum, eigenfunctions, and Koopman modes. In the first low-dimensional example of a classical pendulum system, our method requires significantly fewer dictionary observables than [12, Section 4.3.1, Section 6.3] to compute the Koopman spectrum. The second high-dimensional example on turbulence highlights our method’s ability to detect acoustic vibrations and distinguish the pressure field through Koopman modes. The third example, a real-world high-dimensional neural system, compares NN-ResDMD with three popular methods—Hankel-DMD [25], EDMD with RBF basis, and kernelized-ResDMD [12]—and demonstrates its superiority in identifying and clustering latent dynamic structures. Together, these examples showcase NN-ResDMD’s performance across diverse systems and comprehensively evaluate its capabilities.
Specifically, in all three experiments, we compare NN-ResDMD with Hankel-DMD, a theoretically grounded method that analyzes dynamical systems using time-delayed state measurements (see Appendix A.6.1 for a justification). Although its performance rivals NN-ResDMD in the simple pendulum system, it fails to capture key dynamics in higher-dimensional systems.
@@ 4.1 Pendulum
The pendulum system is a measure-preserving system due to its Hamiltonian nature, which theoretically implies its whole spectrum lies on the unit circle. For its dynamical behaviors, if the initial position of the pendulum is sufficiently far from the peak and the initial angular speed sufficiently small, the pendulum will oscillate; otherwise, the pendulum will pass the peak and rotate. In other words, this complex system exhibits two types of dynamical behaviors: rotation and oscillation. Here we simulate two cases with different numbers of initial points. We choose 90 and 240 initial points uniformly in the domain $[-\pi,\pi]_{per}\times[-15,15]$. Each point evolves 1000 steps with a step size of 0.5. Thus, the total data size in each set is approximately $9\times 10^{4}$ and $2.4\times 10^{5}$, respectively.
As shown in Figure 4, only $N_{K}=300$ observables are needed to approximate the full spectrum, significantly fewer than the nearly 1000 required in [12, Section 4.3.1] for the same data size. Even with a much larger data size (Figure 4), the required observables remain small ($N_{K}=350$), demonstrating the robustness of efficient observables across data sizes.
As shown in Figure 4, we compare our method to four approaches: EDMD, EDMD with Dictionary Learning (EDMD-DL), Hankel-DMD, and ResDMD, on the dataset with 90 initial points to compute the Koopman matrix and its corresponding spectral information. The first three methods (EDMD [11], EDMD-DL [18], and Hankel-DMD are limited to computing eigenvalues associated with the point spectrum. In these experiments, both EDMD and ResDMD use the hyperbolic cross approximation with Hermite functions up to order 15 and Fourier functions up to order 20. Hankel-DMD uses a time delay of 150. While Hankel-DMD yields accurate eigenvalues and shows points near the unit circle which matches the ground truth for this Hamiltonian system, it only captures the point spectrum(eigenvalues) and misses the entire spectrum(eigenvalues + continuous spectrum). It also still suffers from spectral pollution and requires careful tuning of the time delay parameter. With 300 basis functions, ResDMD is still unable to fully capture the entire spectrum, i.e., the unit circle, due to the insufficient number of basis functions. In the original ResDMD work, 964 basis functions using a hyperbolic cross approximation of order 100 were required to adequately cover the spectrum with a dataset of the same size [12, Section 4.3.1]. Our method, in contrast, represents the results using shaded areas that show the pseudospectrum, which is a key feature that can capture the whole spectrum. While the shaded area may appear broad, this actually demonstrates our method’s ability to detect the complete spectrum. This radius of the shaded region accounts for computational uncertainties, as exact spectrum computation is computationally impossible. Theoretically, ResDMD guarantees that as this error tolerance(radius of the shaded region) approaches zero, the pseudospectrum converges to the true spectrum (the unit circle in this case) without spectral pollution. This comparison demonstrates that NN-ResDMD, even with only 300 basis functions, outperforms all four classical methods in terms of capturing the complete spectrum with greater accuracy and fewer basis functions.
NN-ResDMD: $N_{K}=25$
NN-ResDMD: $N_{K}=50$
NN-ResDMD: $N_{K}=100$
NN-ResDMD: $N_{K}=300$
Figure 2: The four plots depict the spectrum of the Koopman operator, constructed using varying dictionary size $N_{K}$ of 25, 50, 100, and 300. Each plot utilizes 90 initial points to illustrate the impact of increasing the dictionary size on approximating the spectrum of the Koopman operator.
NN-ResDMD: $N_{K}=50$
NN-ResDMD: $N_{K}=150$
NN-ResDMD: $N_{K}=250$
NN-ResDMD: $N_{K}=350$
Figure 3: Same example as Figure 4 but with larger data size, using 240 initial points to show the effect of increasing dictionary size on approximating the Koopman operator spectrum.
@@
Figure 4: Comparison with classical methods. The four plots above represent the spectral information obtained from a $300\times 300$ Koopman matrix, calculated using four methods: EDMD, EDMD with Dictionary Learning (EDMD-DL), Hankel-DMD, and ResDMD. The illustrated eigenvalue spectra of the Koopman operator highlight the differences in results produced by these methods.
@@ 4.2 Turbulence
Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics, where Kernel ResDMD has been particularly successful in capturing such patterns and detecting acoustic vibrations in the turbulence system. However, Kernel ResDMD requires careful selection of kernel functions, while NN-ResDMD bypasses this by using neural networks to train observables and compute Koopman modes.
We demonstrate this by applying NN-ResDMD to the turbulence system using the dataset from [12, Section 6.3]. The ground truth in the first plot of Figure 5 represents a high-dimensional pressure field distribution (approximately 30,000 spatial dimensions) around an airfoil, with a clear distinction between the upper and lower surfaces. Technically, we apply truncated Singular Value Decomposition (SVD), select 300 observables, compute Koopman modes, and project them back into the original state space.
In Figure 5, the first Koopman mode estimated by NN-ResDMD with the smallest residual value successfully highlights a clear global spatial separation that aligns with the pattern observed in the original pressure field. This advantage allows the first Koopman mode to directly distinguish spatial features present in the true pressure field, making it a powerful tool for interpreting complex fluid dynamics data. Subsequent Koopman modes also reveal strong acoustic waves that are critical in various aeronautical engineering fields. In contrast, Kernel ResDMD with a generic normalized Gaussian kernel function, as shown in the original work, is unable to produce a Koopman mode similar to the first Koopman mode from NN-ResDMD that clearly distinguishes the pressure field. For comparison, we also plot four Koopman modes computed by Hankel-DMD with a time delay of 5, corresponding to the four smallest residual values, which similarly do not reveal the pressure field patterns as in NN-ResDMD. These results are presented in Appendix Figure 7 and Appenxic A.6.2.
Figure 5: The plots illustrate turbulence detection with Koopman modes computed by 300 observables. The first plot shows a 2D scatter plot of the pressure field, while the other plots display various Koopman modes, each labeled with corresponding residuals. The small residual values in the figures associated with the Koopman modes confirm the estimation accuracy.
4.3 Identification of neural dynamics in mice visual cortex

### v3

@@ 3.2 Residual-based Koopman-learning Network (ResKoopNet)
where each column of the matrix $V$ is an (right) eigenvector $\mathbf{v}_{i}$ of the matrix $\tilde{K}$. Thus, with dictionary $\mathbf{\Psi}$ fixed, we can obtain a closed form for the optimal Koopman matrix $\tilde{K}$ as following
$\tilde{K}=G^{\dagger}A,$ (3.6)
where $G=\frac{1}{m}\Psi_{X}^{*}\Psi_{X},A=\frac{1}{m}\Psi_{X}^{*}\Psi_{Y}$ (See A.3 for more details on deriving the Eq. (3.5) and Eq. (3.6)).
Although Eq. (3.6) resembles the EDMD formulation for computing the Koopman matrix representation $\tilde{K}$, ResKoopNet and ResDMD fundamentally differ from EDMD in their theoretical foundation. While they share a similar matrix expression, this expression is derived from minimizing the spectral residual in Eq. (3.3) rather than the prediction error that EDMD minimizes, as shown in Appendix A.3. Specifically, our approach explicitly incorporates $\mathcal{K}^{*}\mathcal{K}$ as shown in Eq. (3.2), focusing on spectral accuracy rather than observable prediction. This distinction leads to significantly different performance in capturing the spectral properties of the Koopman operator.
Remark 3.2.
To ensure the numerical stability when computing $G^{\dagger}$, we add a small perturbation, i.e., $\tilde{K}=(G+\sigma I)^{-1}A$ for some small number $\sigma>0$.
As shown in Eq. (3.6), ResKoopNet provides an explicit expression for the optimal Koopman matrix $\tilde{K}$ given a fixed dictionary function $\mathbf{\Psi}$. Although the formulation is similar to EDMD, the key distinction between ResKoopNet and EDMD lies in their loss functions; more specifically, while EDMD minimizes the overall prediction error $\|\Psi_{Y}-\Psi_{X}\tilde{K}\|_{F}^{2}$, ResKoopNet minimizes the residual projected onto the estimated eigenspace, i.e., $\|(\Psi_{Y}-\Psi_{X}\tilde{K})V\|_{F}^{2}$. This eigenspace projection fundamentally changes the optimization landscape by weighting errors along spectral directions, ensuring that the learned Koopman matrix better captures the underlying eigenstructure rather than merely minimizing average prediction error. Therefore, they arise from distinct theoretical foundation.
From Residual to Machine Learning This section explains how a simple neural network is integrated into ResKoopNet. In the optimizer as shown in Figure 1, the neural network parameterizes the dictionary $\mathbf{\Psi}(x;\theta)$ and minimize the parametrized total spectral residual $J(\theta)$ given in the following:
$J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2},$ (3.7)
@@
where $K(\theta)$ and $V(\theta)$ depend on the parameters $\theta$. The optimal Koopman matrix $\tilde{K}(\theta)$ is computed following Eq. (3.6):
$\tilde{K}(\theta)=G(\theta)^{\dagger}A(\theta).$ (3.8)
@@
The algorithm alternates between updating $K(\theta)$ via Eq. (3.8) and optimizing $J(\theta)$ w.r.t. $\theta$ by stochastic gradient descent via Eq. (3.7). While it is possible to optimize both $K(\theta)$ and $J(\theta)$ simultaneously, as done in Takeishi et al. [35] and Otto and Rowley [30], our separate procedure ensures computational efficiency and numerical stability [23].
@@
Computing Algorithm In our neural network implementation, we include some non-trainable functions to enhance the dictionary computation. Specifically, we add a vector of constant one and the coordinates of the state space as non-trainable basis in the output layer, which helps avoid trivial solutions, i.e., $J=0$ for some initial $\theta$. For the network architecture, we build a simple feedforward neural network with three hidden layers, where each hidden layer size can be specified during training. We use the hyperbolic tangent (tanh) function as the activation function for the hidden layers and employ the Adam optimizer for updating the network parameters $\theta$. Adam is particularly well-suited for this task due to its ability to adapt the learning rate for each parameter, which can lead to faster convergence in the alternating optimization process between the network parameters $\theta$ and the Koopman matrix $K(\theta)$ formulation. The computing steps are illustrated in the following Algorithm 1.
@@
1 Initialize Initialize $\theta$, thus initializing $\mathbf{\Psi}(\theta)$ ;
@@ 6       Update $\tilde{K}(\theta)=(G(\theta)+\sigma I)^{-1}A(\theta)$ and $V(\theta)$ ;
Output: $\tilde{K}(\theta)$, eigenpairs $\{(\lambda_{i},\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i})\}_{i=1}^{N_{K}}$ and pseudospectrum $\{z_{j}:\tau_{j}<\varepsilon\}$ (See Eq. (3.4)).
@@
ResKoopNet’s practical advantages come with computational demands due to its iterative optimization. Each gradient update scales linearly with system dimensionality and network parameters. While individual steps are lightweight, repeated least-squares optimizations make convergence slower than standard numerical methods, with stochastic gradient descent achieving an $O(1/n)$ rate. We also give a brief discussion on the convergence analysis of ResKoopNet in Appendix A.4, which leverages existing results from approximation theory in the Barron space [14, 42].
@@ 4 Application in physical and biological systems
In this section, we present three examples to demonstrate ResKoopNet’s effectiveness in estimating key components of Koopman Mode Decomposition: spectrum, eigenfunctions, and Koopman modes. In the pendulum example, we compare with EDMD, EDMD-DL, Hankel-DMD, and ResDMD, and show that ResKoopNet captures the full Koopman spectrum with significantly fewer dictionary observables than required by ResDMD, while other methods either produce spectral pollution or miss the continuous spectrum. Our turbulence analysis demonstrates ResKoopNet’s ability to detect not only acoustic vibration and turbulent fluctuation (as in ResDMD [8, Section 4.3.1, Section 6.3]) but also to recover the fundamental pressure field structure through the lowest-residual Koopman mode, which is a feature not captured by Hankel-DMD using the same dataset. In the neural dynamics example, we compare against three methods across five mice datasets, with ResKoopNet achieving superior clustering of neural states corresponding to different visual stimuli, as quantified by the Davies-Bouldin index. Note that we include Hankel-DMD [1] as a consistent benchmark across all experiments due to its theoretical guarantees and applicability to both low and high-dimensional systems without requiring predefined dictionaries. It uses time-delayed state measurements by Takens Embedding Theorem [36] instead of using a pre-defined dictionary or applying Galerkin approximation as in EDMD and ResDMD methods (see Appendix A.8.1 for more details.)
@@ 4.1 Pendulum
The pendulum system is a measure-preserving Hamiltonian system, which theoretically constrains its spectrum to the unit circle. The system is governed by:
$\ddot{\theta}=\sin(\theta),\quad(\theta,\dot{\theta})\in[-\pi,\pi]_{\text{per}}\times[-15,15],$
where $\theta$ represents the angular displacement from equilibrium and $\dot{\theta}$ the angular velocity. The dynamics exhibit two distinct regimes: oscillatory motion for low-energy initial conditions, and rotational motion when the initial energy exceeds the separatrix threshold [24].
For our analysis, we sample 90 and 240 initial conditions uniformly distributed in the domain, each evolving for 1000 time steps with $\Delta t=0.5$. This gives datasets of approximately $9\times 10^{4}$ and $2.4\times 10^{5}$ snapshots, respectively. We choose 3 hidden layers and use 300 neurons and 350 neuron in each hidden layer in both cases, respectively.
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
Figure 2: The four plots depict the spectrum of the Koopman operator, constructed using varying dictionary sizes $N_{K}$ of 25, 50, 100, and 300. Each plot utilizes 90 initial points to illustrate the impact of increasing the dictionary size on approximating the spectrum of the Koopman operator.
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
Figure 3: Same example as Figure 2 but with larger data size, using 240 initial points to show the effect of increasing dictionary size on the Koopman spectrum approximation.
@@
Figure 4: Comparison with classical methods. Eigenvalue spectra computed from a $300\times 300$ Koopman matrix using EDMD, EDMD-DL, Hankel-DMD, and ResDMD.
As shown in Figure 2, ResKoopNet requires only $N_{K}=300$ observables to approximate the full spectrum. While the original ResDMD work reported using 964 basis functions [8, Section 4.3.1], our additional tests indicate that approximately 460 basis functions can produce reasonable circular approximations with ResDMD (see Figure 7 in Appendix A.7.1). Nevertheless, ResKoopNet still achieves comparable accuracy with significantly fewer observables. This efficiency persists even with larger sizes of data points (Figure 3), where $N_{K}=350$ observables remain sufficient. Figure 4 compares ResKoopNet with four classical methods using the 90-initial-point dataset. Both EDMD and ResDMD employ Hermite functions (order 15) for $\theta$ and Fourier functions (order 20) for $\dot{\theta}$, while Hankel-DMD uses a time delay of 150. Though Hankel-DMD produces eigenvalues near the unit circle, it only captures the discrete point spectrum and cannot reconstruct the continuous spectral components. We include Hankel-DMD here for consistency in benchmarking across all experiments, despite its limitations in higher-dimensional systems examined later. The shaded region in the ResDMD plot represents the $\epsilon$-pseudospectrum, which approximates regions where the spectral residual is smaller than a threshold value; however, with 300 basis functions, ResDMD fails to fully recover the unit circle. Overall, this pendulum example demonstrates that ResKoopNet achieves accurate and complete spectral recovery using much fewer observables than other classical methods.
@@ 4.2 Turbulence
Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics [33, 28]. Figure 5(a) shows the ground truth pressure distribution, where a clear asymmetry exists between the upper and lower airfoil surfaces. This fundamental spatial pattern is critical for analyzing aerodynamic performance and structural integrity. While the original Kernel-ResDMD method has successfully detected acoustic vibration and turbulent fluctuation shown in Colbrook and Townsend [8, Section 6.3], it fails to recover the global spatial structure of the pressure field. In contrast, ResKoopNet achieves a significant improvement by accurately reconstructing this dominant spatial pressure field structure itself from its leading Koopman mode as shown in Figure 5(b), which contains essential information about the flow characteristics around the airfoil.
The turbulent flow dataset from Colbrook and Townsend [8, Section 6.3] models a two-dimensional airfoil system with Reynolds number $3.88\times 10^{5}$ and Mach number 0.07. The data captures a pressure field at 295,122 spatial points across 798 time steps, sampled every $2\times 10^{-5}$ seconds. Given the extreme dimensionality of this $X\in\mathbb{R}^{798\times 295,122}$ matrix, we apply truncated SVD on the data matrix to retain only the top 150 components and project the data via $XV=US\in\mathbb{R}^{798\times 150}$. After computing Koopman modes in this reduced state space using ResKoopNet, we map them back to the original dimensions through multiplication with $V$. To maintain consistency with the baseline, we use the same number of observables $N_{K}=250$ as in Colbrook and Townsend [8, Section 6.3], including 99 trainable neural network-based functions, one constant function, and 150 principal components obtained from the truncated SVD.
Figure 5(b) shows the first Koopman mode computed by ResKoopNet, which corresponds to the mode whose eigenvalue has the largest absolute real part, and it also has the smallest spectral residual among all computed modes. This mode not only captures the most dominant dynamics in the system but also closely resembles the original pressure field. This correspondence demonstrates that ResKoopNet is able to recover the principal spatial feature of the turbulent flow, something that Kernel-ResDMD and Hankel-DMD could not achieve, which also justifies that the proposed method ResKoopNet can effectively address the spectral inclusion problem where traditional methods often fail to recover the complete spectrum of the Koopman operator. This mode likely corresponds to the largest singular value observed in the SVD, as shown in Figure 5(e) and (f). The dominant singular value highlights the strongest spatial mode in the data, which aligns with the physical air pressure field and its global structure.
In addition to the first mode, Figure 5(c) and (d) show the third and seventh Koopman modes computed by ResKoopNet, associated with the eigenvalues having the third and seventh largest absolute values of their real parts. These two modes have the second and third smallest spectral residuals and correspond to acoustic vibration and turbulent fluctuation, as also observed in the Kernel-ResDMD study [8, Section 6.3]. The ordering of the Koopman modes in Figure 5(b)–(d) is based on the ascending values of spectral residual, indicating how well each mode aligns with the spectrum of the underlying Koopman operator.
For further comparison, we include in Appendix A.8.2 four Koopman modes computed using Hankel-DMD with a time delay of 5, selected similarly based on the smallest spectral residual. These modes, however, do not reveal the global pressure pattern observed in ResKoopNet’s first mode, which further validates the unique capability of ResKoopNet to recover large-scale coherent structures in turbulent systems.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
(e) (f)
Figure 5: Turbulence detection using Koopman modes from 250 observables. (a) shows the original 2D pressure field. (b) shows Koopman mode 1 which has the smallest residual and closely matches the original pressure field. (c) and (d) show the acoustic vibration and turbulent fluctuation by the Koopman modes with the 2nd and 3rd smallest spectral residual. (e) and (f) show the plot of all singular values and first 150 singular values when applying truncated SVD method, respectively.
4.3 Neural dynamics identification in mice visual cortex
Figure 6: ResKoopNet outperforms Hankel-DMD in identifying latent dynamic structures in neural signals with a dictionary size of 50. (A) (Top) 50 Koopman eigenfunctions estimated by ResKoopNet across 6 states characterized by different video stimuli in an example mouse. Each trial contains 300 data points (10s at 50Hz). (Bottom) 50 Koopman eigenfunctions approximated by Hankel-DMD, each 50 points long, reflecting the dimension of the Hankel matrix. (B) 2D representation of Koopman eigenfunctions for all tested mice, computed by ResKoopNet and reduced via Multidimensional Scaling (MDS). Trials of the same state cluster well. (C) Same as (B) but computed with Hankel-DMD, showing no clear state separation. (D) 2D representation of Koopman eigenfunctions for the first mouse, computed by EDMD with an RBF basis. See Appendix Figure 14 for full results. (E) Same as (D) but computed with Kernel-ResDMD. See Appendix Figure 15 for full results. (F) Davies-Bouldin Indices (DBIs) evaluating clustering quality across four methods (ResKoopNet, Hankel-DMD, EDMD+RBF, and Kernel-ResDMD) for five mice. Lower DBI values for ResKoopNet indicate better clustering.